### v1

# SourceNet: Interpretable Sim-to-Real Inference on Variable-Geometry Sensor Arrays for Earthquake Source Inversion

Abstract
Inferring high-dimensional physical states from sparse, ad-hoc sensor arrays is a fundamental challenge across AI for Science, as they are complicated by irregular geometries and the profound Sim-to-Real gap in physical modeling. Taking earthquake source characterization as a representative challenge, we address limitations in conventional deep learning: CNNs demand fixed grids, while pooling-based architectures (e.g., DeepSets) struggle to capture the relational wave physics. Here, we propose SourceNet, a Transformer-based framework that treats the sensor array as a flexible set to model arbitrary geometries. To bridge the reality gap, we introduce Physics-Structured Domain Randomization (PSDR). Instead of forcing feature alignment, PSDR randomizes the governing physical dynamics by varying velocity structures, propagation effects, and sensor availability, to force the model to learn robust representations invariant to unmodeled environmental heterogeneity. By pre-training on 100,000 synthetic events and fine-tuning on $\sim$2,000 real world events, SourceNet achieves state-of-the-art precision on held-out real data. This demonstrates exceptional data efficiency, and matches classical solvers while enabling real-time processing. Remarkably, interpretability analysis reveals that the model shows scientific-agent-like features: it autonomously discovers geometric information bottlenecks and learns an attention policy that prioritizes sparse sensor placements, effectively recovering principles of optimal experimental design from data alone.
AI for Science, Inverse Problems, Sim-to-Real, Geophysics, Set Transformer
1 Introduction
How do we reconstruct a complex physical event, whether it is a distant earthquake, a cosmic radio burst, or a climate anomaly, using only a handful of scattered sensors? This is the archetypal challenge across the physical sciences (Bergen et al., 2019; Bouman et al., 2016; Ravuri et al., 2021): inferring complex physical source parameters $\theta$ from a sparse set of indirect, noisy observations $X=\{x_{1},\dots,x_{n}\}$. Unlike images or video where data pixels are fixed on a regular grid, scientific data is often collected by ad-hoc sensor arrays where the number of sensors varies per event and their spatial arrangement is irregular.
While classical inverse solvers (Tarantola, 2005), which typically rely on iterative optimization on pre-computed Green’s functions, are theoretically robust, they are computationally costly for real-time applications. Deep learning offers a compelling alternative via amortized inference (Cranmer et al., 2020) where the inverse operator is trained once and applied in various inferences with instantaneous prediction. However, there is a fundamental geometric mismatch between standard AI architectures and physical data. Convolutional Neural Networks (CNNs) (LeCun et al., 2015) need rigid grids that force researchers to use artificial interpolation. Graph Neural Networks (GNNs) can handle irregularity but often require fixed topologies (Bronstein et al., 2017). To address these challenges, we need permutation-invariant set architectures capable of modeling real-world sensor data with flexible geometries (Zaheer et al., 2017; Qi et al., 2017).
Compounding the geometric challenge is the Sim-to-Real gap (Tobin et al., 2017). In science disciplines like geophysics and robotics, obtaining ground-truth labels for real-world observations is expensive and labor-intensive, therefore models have to trained on physics-based simulations, which inevitably lack the dirty reality including instrumental noise, realistic wave propagation effect, and structural heterogeneity (Karniadakis et al., 2021). Hence, naive training on idealized simulations faces a substantial distribution shift and is difficult to generalize to various real-world data. A trustworthy AI for Science framework must learn invariant physical operations that hold true despite these distribution shifts.
Here, we present SourceNet, an end-to-end, interpretable framework for solving inverse problems on variable sensor sets. We take earthquake focal mechanism inversion as our task. We have three methodological advancements:
  1. 1.
    We bridge the Sim-to-Real gap not by feature alignment, but by randomizing the physics itself (e.g., velocity models, sensor dropouts), forcing the model to learn representations invariant to crustal heterogeneity.
  2. 2.
    We demonstrate that self-attention is theoretically superior to pooling-based baselines (e.g., DeepSets) for wave physics. By modeling pairwise dependencies, SourceNet effectively aggregates full-waveform information including scattered coda phases typically ignored by classical solvers, to achieve SOTA accuracy on both synthetic benchmarks and real-world catalogs, even under sparse station coverage.
  3. 3.
    Finally, through Explainable AI (XAI) analysis, we reveal that SourceNet acts as a scientific agent, as it autonomously identifies the geometric information bottleneck of the sensor network and learns a global policy that prioritizes sparsely distributed sensor orientations. This effectively maximizes information gain without explicit supervision.
2 Related Work
2.1 Deep Learning on Sets and Irregular Grids
Handling varying cardinality inputs require permutation-invariant architectures. Early set-based architectures, such as DeepSets (Zaheer et al., 2017) and PointNet (Qi et al., 2017), can achieve permutation invariance by applying independent encoders followed by symmetric global pooling. While effective for simple classification tasks, we argue these are insufficient for wave physics because pooling reduces the set to a sum, which limits their ability to model relational context (e.g., amplitude differences between stations) critical for inversion. Graph Neural Networks (GNNs) (Kipf, 2016; Battaglia et al., 2018) model dependencies explicitly but typically assume fixed edges or require expensive dynamic graph construction ($O(N^{2})$) (Wang et al., 2019). More recently, Neural Operators (e.g., DeepONet(Lu et al., 2021), FNO(Li et al., 2020)) have emerged to learn mappings between infinite-dimensional function spaces. However, these methods often assume dense sampling of fixed control points. In contrast, Set Transformers (Lee et al., 2019) leverage self-attention to model pairwise interactions dynamically. In the context of wave physics, attention-based mechanisms offer a theoretical advantage over pooling-based baselines by preserving radiation patterns on sparse arrays.
2.2 Sim-to-Real Transfer and Domain Randomization
Bridging the simulation-to-reality gap is a key challenge of domains ranging from robotics (Tobin et al., 2017) and neurobioscience (Gonçalves et al., 2020) to geophysics. In these fields, simulations provide abundant data but inevitably simplify the complex, unmodeled dynamics of the physical world (Pan & Yang, 2009). Traditional unsupervised domain adaptation methods aim to align feature distributions via adversarial training (Ganin et al., 2016) or minimizing maximum mean discrepancy (Long et al., 2015), but these require direct access to the target domain during training and can be unstable. Domain randomization (DR) (Tobin et al., 2017) takes a different approach: it randomizes simulator parameters to force the model to learn invariants (Peng et al., 2018). While DR has become standard in robotics for handling visual and contact dynamics (Akkaya et al., 2019), scientific inverse problems require a deeper form of adaptation because the uncertainty lies in the physics themselves. Our SourceNet extends DR by randomizing the governing physical processes (velocity structure, wave propagation effect, scattering) rather than just visual appearance. This not only solves the specific challenge of seismic inversion but offers a generalizable blueprint for learning robust operators in many physical system with imperfect simulations.
2.3 Machine Learning for Geophysical Inverse Problems
Deep learning has substantially transformed observational seismology, particularly in discriminative tasks where models like PhaseNet (Zhu & Beroza, 2019) and Transformer-based pickers (Mousavi et al., 2020) have scaled earthquake catalogs to millions by automating phase association and denoising (McBrearty & Beroza, 2023; Ross et al., 2019). However, end-to-end parameter inversion remains an open challenge. Unlike detection tasks, inversion requires capturing subtle radiation patterns across irregular, variable-cardinality arrays. Standard CNN-based solvers (e.g., FMNet (Kuang et al., 2021)) typically impose artificial grids which limits their generalization to real-world ad-hoc networks. While Graph Neural Networks (GNNs) (Zhang et al., 2022) and shift-and-stack operators (Zhu et al., 2022) address geometric irregularity, they often incur quadratic computational costs, or implicitly treat sensors as independent evidence and ignore the directional dependencies between stations. Recent Transformers like TEAM (Münchmeyer et al., 2021) treat sensors as tokens yet rely on discrete positional encodings, which lacks the continuous spatial resolution required for precise physical inference.
2.4 Interpretability as Emergent Strategies in Science
Building trust in scientific AI requires more than just accuracy. Standard visualization tools (e.g., Grad-CAM (Selvaraju et al., 2017)) typically show where a model looks, but may not explain why specific data points matter physically (Adebayo et al., 2018; Rudin, 2019). Recent work in mechanistic interpretability attempts to bridge this gap by verifying whether models have internalized abstract physical concepts, such as laws or symmetries, rather than simply memorizing dataset artifacts (Cranmer et al., 2020; Iten et al., 2020). We address this interpretability challenge from the perspective of Optimal Experimental Design (OED) (Krause et al., 2008; Chaloner & Verdinelli, 1995), a statistical approach dedicated to organizing observations to maximize information gain (thus reducing uncertainty). While modern approaches have begun to approximate OED using variational deep learning (Foster et al., 2019), they typically treat it as a separate optimization loop. We hypothesize that OED strategies can be emergent, and an attention-based model, when trained to solve a hard inverse problem, could implicitly discover these optimal sensing policies. By analyzing the Transformer’s attention weights, we aim to show that the model autonomously learns to prioritize sensors based on their information value, which effectively recovers experimental design principle purely from data.
3 Problem Formulation and Methodology
We formalize the earthquake source characterization as a regression task mapping a variable set sensor data to a physical source mechanism.
Figure 1: The SourceNet Architecture. A hierarchical Set Transformer designed for irregular sensor arrays. The model operates in three stages: (1) Multi-modal Station Encoders process P/S-waves (time & spectral domains) and scalar metadata into physics-aware embeddings. (2) Self-Attention models global pairwise interactions to resolve local ambiguities. (3) Attention Pooling dynamically weights stations based on information gain to regress the final moment tensor.
The physical source ($y$). Our target is the seismic event $\mathcal{E}$ parameterized by a latent physical state $\mathbf{y}\in\mathbb{R}^{6}$. This state encapsulates the Moment Tensor (3*3 symmetric tensor) (Aki & Richards, 2002), representing the fault orientation and slip direction (5 independent zero-trace $M_{ij}$ components), and the moment magnitude $M_{w}$ indicating the energy scale (Kanamori, 1977).
The Observation Set ($X$). Unlike images defined on regular grids, a seismic network is an ad-hoc collection of sensors. We represent the observation as an unordered set $\mathbf{X}=\{(\mathbf{w}_{i},\mathbf{s}_{i})\}_{i=1}^{N}$, where $N$ is the number of stations (varies per event). Each element in the set is a multi-modal tuple: $\mathbf{w}_{i}\in\mathbb{R}^{C\times T}$ represents the dual-domain waveform features, including both the raw velocity time-series (for phase information) and their spectral representations (for frequency-dependent properties); $\mathbf{s}_{i}\in\mathbb{R}^{d}$ represents scalar metadata including station azimuth, epicentral distance, locations, and the amplitude and amplitude ratios of the waveforms. Thus, they encode the physical attenuation from distance-based amplitude decay necessary to unravel the source magnitude.
Learning Objectives. We aim to learn an inverse operator $f_{\theta}:\mathbf{X}\to\mathbf{y}$ that maps the observed sensor set to the source parameters. This function must satisfy permutation invariance, i.e., $f_{\theta}(\pi(\mathbf{X}))=f_{\theta}(\mathbf{X})$ for any permutation $\pi$ (Zaheer et al., 2017). Since real-world labels are scarce, we train primarily on a synthetic source domain $\mathcal{D}_{\text{Sim}}$ but optimize for performance on the real-world target domain $\mathcal{D}_{\text{Real}}$. Our strategy is to inject physical priors into $\mathcal{D}_{\text{Sim}}$ such that the model learns invariant operators robust to the distribution shift.
3.1 Physics-Structured Domain Randomization (PSDR)
Bridging the reality gap in physical sciences requires more than just adding Gaussian noise. Real-world seismic data is distorted by the unknown 3-D earth structures, complex scattering, and site and sensor responses that standard simulations cannot capture. To ensure our model generalizes to real data without ever seeing them during training, we propose Physics-Structured Domain Randomization (PSDR).
Figure 2: Physics-Structured Training Curriculum. We bridge the reality gap using a two-stage protocol. First, we train on synthetics augmented with aleatoric physical uncertainties (e.g., randomized velocity models, real noise injection, and sensor dropouts) to enforce invariant feature learning. We then fine-tune on a smaller set of real events using a weighted random sampler to correct for the natural imbalance of faulting mechanisms.
The core idea is simple: instead of unstable domain adaptation to align real data with synthetic distributions, we make our simulations as chaotic and diverse as the real world. We define our data generation process as:
$\mathbf{X}_{\text{aug}}=\mathcal{M}\left(\mathcal{T}(S(\mathbf{y},\phi))+\mathbf{n}\right)$ (1)
where $S$ is the physics-based wavefield simulator. We introduce four specific types of randomization to mimic real-world entropy:
  • Earth Structure ($\phi$): Standard simulations assume a single, perfect 1D Earth model. The real crust, however, is heterogeneous. To prevent the model from memorizing a specific velocity structure, we sample from a library of 17 distinct 1D velocity models derived from CRUST1.0 (Laske et al., 2013) for the Southern California region during synthetic generation. This forces the network to learn wave features that are invariant to the propagation medium.
  • Realistic Noise ($\mathbf{n}$): Gaussian noise cannot capture the complex spectral properties of Earth’s background vibration (e.g., ocean microseisms or traffic). We strictly sample noise vectors $\mathbf{n}$ from a library of real-world ambient recordings and superimpose them onto synthetics. This teaches the model to separate signal from realistic noise sources.
  • Signal Distortion ($\mathcal{T}$): Real waveforms are different from theoretical predictions due to unmodeled 3D scattering and site effects (Sato et al., 2012). We apply stochastic time shifts and amplifications, and we also superimpose simulated scattering coda (exponentially decaying tails) following the body waves. This serves as a regularization to prevent the model from overfitting to the unrealistically clean 1D Green’s functions and promote using the full wavetrain data.
  • Network Availability ($\mathcal{M}$): Real sensors can fail or have gaps. We apply a stochastic masking operator $\mathcal{M}$ (Bernoulli dropout) to randomly discard stations during training, to make the model robust to sparse and variable geometries.
Validation via Manifold Alignment. To validate our physics-structured approach, we compare the latent spaces of the baseline model trained on clean synthetics and SourceNet trained on PSDR, where both models are fine-tuned on the same subset of real data. As shown in the t-SNE (Maaten & Hinton, 2008) visualizations in Figure 3, standard fine-tuning is insufficient to bridge the reality gap, shown by the disjoint clusters of the naive baseline model, and it memorize real samples rather than learning generalized physics. In contrast, SourceNet achieves perfect manifold alignment. This proves that PSDR is a prerequisite for effective transfer learning, allowing the model to learn domain-invariant physical features that fine-tuning alone cannot recover.
3.2 SourceNet Architecture
SourceNet is designed to mirror how a seismologist would analyzes an event: first extracting features from each station (phases, amplitudes, location), and then synthesizing this information globally to constrain the source. As illustrated in Figure 1, the model consists of three functional modules:
1. Domain-Specific Station Encoders. Generic time-series models usually treat all inputs identically, but time series data like seismic waves can contain distinct physical components. Therefore, we decompose the input for the $i$-th station into three parallel streams, thus creating a multi-modal view of the data:
  • P-Wave Tower: We extract a 6-second window ([1s,+5s][-1s,+5s] relative to the P wave onset) filtered in the low-frequency band (0.1–2 Hz) to remove the drift and noise. This stream processes both the raw velocity waveforms and their spectral representations via a 1D ResNet (He et al., 2016). This stream captures not only the high-frequency onset polarity, but also the early crustal phases in the coda. As our analysis reveals, these converted phases contain rich constraints often ignored by traditional first-motion solvers.
  • S-Wave Tower: We also extract the S-wave window (same filtering and duration). Unlike the P-wave stream, this tower specializes in capturing shear energy which is physically distinct from and orthogonal to the P-wave particle motion. This allows the model to disentangle shear from compressional features and maximize the constraints on the source mechanism.
  • Scalar Tower: Standard Transformers (Vaswani et al., 2017) use discrete positional encodings (like integers in a sentence), which can fail for irregular sensor arrays. We instead project a vector of 20 explicit physical features through an MLP. This vector includes the geometric metadata: Station coordinates (lat/lon), azimuth, epicentral distance, and source depth; the amplitude information: Max absolute amplitudes for both time-domain and spectral channels (6 for P-wave, 6 for S-wave); and 3 P/S amplitude ratios. This generates a continuous embedding that allows the scale-invariant Transformer to recover the absolute Moment Magnitude ($M_{w}$) and the source radiation patterns.
These three views are further concatenated to form a rich station embedding $h_{i}$, transforming the raw data into a physics-aware feature set.
Figure 3: Validation of Physics-Structured Domain Randomization. (a) Comparison of clean synthetic data (top) versus PSDR-augmented data (bottom), which incorporate realistic scattering and noise. (b) t-SNE visualization of the feature space. A naive baseline (Left) shows disjoint clusters for synthetic and real data, indicating a failure to generalize. In contrast, SourceNet (Right) achieves good manifold alignment, which shows that the model has learned invariant physical operators robust to the distribution shift.
2. Global Context via Self-Attention. A challenge to the source inversion is that local observations are ambiguous: a single sensor cannot distinguish between different physical parameters without reference to other sensor data. Standard pooling (DeepSets) can fail because it aggregates these ambiguous features directly and masks the relational context. We employ a Transformer Encoder to model these relations. Through self-attention, each station queries the entire array, aggregating information from distant sensors to update its own representation. This process contextualizes the local data features against the global radiation pattern, which allows each station’s embedding to encode not just what it saw, but how it fits into the collective source mechanism.
3. Weighted Aggregation (Attention Pooling). Real-world inference requires adaptability since the key data is not fixed but varies significantly from event to event depending on the source location and radiation pattern. Instead of a static average such as average pooling, we use attention pooling (Ilse et al., 2018) for event-specific aggregation. This approach dynamically weights which subset of sensor data provides the key constraints for an event in query, thereby optimizing the global representation specific to the corresponding focal mechanism solution.
4 Experimental Setup
We design our experiments to rigorously test SourceNet’s ability to learn from physically diverse simulations and transfer to a biased, real-world seismic catalog. Our study focuses on the Southern California region, leveraging its dense seismic network and well-characterized earthquake catalog (Yang et al., 2012).
4.1 Datasets
Synthetic Pre-training Domain. We generate a large-scale dataset of 100,000 synthetic events. To ensure the model learns a general physical operator, we introduce two key sources of diversity:
  1. 1.
    Uniform Source Mechanism Sampling: Real-world catalogs are often dominated by one type of faulting (e.g., strike-slip in California). To mitigate this bias, we sample focal mechanisms uniformly from the space of all possible double-couple sources. We parameterize this using the 6-component moment tensor ($M$), which provides a more continuous and uniform sampling space than the non-unique strike-dip-rake angles.
  2. 2.
    Realistic Geometries: The event hypocenters and the available station layouts are sampled directly from the real Southern California catalog (Yang et al., 2012). To simulate variable network density, for events with more than 50 available stations, we randomly sub-sample a set of 30-50 stations for each training instance.
Real-world Fine-tuning Domain. We use a high-quality catalog of 2,544 $M>3.0$ earthquakes in Southern California with focal mechanism solutions from (Yang et al., 2012). These events are selected for their reliability (Quality A & B) and acceptable station coverage (¿5 stations) (Hardebeck & Shearer, 2002).
4.2 Training Protocol
Our training follows a two-stage curriculum:
  1. 1.
    Pre-training on Synthetics: We pre-train SourceNet for 150 epochs on the synthetic dataset using an AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of $2*10^{-4}$. The data is split into 80% training, 10% validation, and 10% testing.
  2. 2.
    Fine-tuning on Real Data: We then fine-tune the entire network on the real-world dataset. Due to the strong prevalence of strike-slip events, a naive fine-tuning would cause the model to forget rare mechanisms. To prevent this, we employ a weighted random sampler. We discretize the 5-D moment tensor space into bins and assign higher sampling weights to events from less populated bins (e.g., thrust or normal faulting). This ensures the model receives balanced exposure to all faulting types. The fine-tuning runs with a low learning rate of $2*10^{-6}$, with a 70% training, 15% validation, and 15% testing split.
5 Experimental Results and Analysis
We evaluate SourceNet on both the held-out synthetic test set (for theoretical upper bound) and the real-world Southern California catalog (to measure Sim-to-Real transfer).
5.1 Sim-to-Real Generalization
As shown in Figure 4 (Top), SourceNet predicts source parameters very well on the synthetic test set, with Mean Absolute Errors (MAE) for moment tensor ($M_{ij}$) ranging from $0.05$ to $0.08$. This shows that the architecture can well approximate the nonlinear wave physics. When applied to real-world data (Figure 4, Bottom), we observe a degradation due to the inevitable reality gap. However, the predictions remain nicely correlated with ground truth, and the magnitude prediction remains robust ($MAE=0.11$) although real-world waveforms are distorted by scattering. As we demonstrate in Section 5.3, this robustness is directly attributable to our multi-modal design, which disentangles geometric spreading (via the scalar tower) from source physics.
Figure 4: Sim-to-Real Inference Performance. Evaluation on the held-out real-world catalog. (a) Predicted vs. true Magnitude ($M_{w}$) demonstrates high accuracy ($MAE=0.11$), confirming that the scalar tower successfully disentangles source energy from path attenuation. (b-c) Normalized Moment Tensor components ($M_{xx},M_{xy}$) maintain robust linearity despite the domain gap. (d) The Kagan angle error distribution (Mean $26.18^{\circ}$, Median $19.37^{\circ}$) confirms that SourceNet achieves SOTA precision and effectively hits the label uncertainty floor of the manual catalogs.
5.2 Benchmarking against SOTA
We benchmark SourceNet against established baselines: the authoritative Southern California catalog (Yang et al., 2012) and recent deep learning solvers. On our operational test set that mixes both good (quality A) and acceptable (quality B) events, SourceNet achieves a mean Kagan angle of $26.18^{\circ}$ and a median of $19.37^{\circ}$. Given that ground-truth uncertainties for Quality B events range up to $35^{\circ}$ (Yang et al., 2012), our error effectively hits the label noise floor. This suggests the model is not merely fitting noisy labels but has learned a robust physical operator more consistent than the noisy manual labels used for supervision.
Comparisons with recent deep learning benchmarks highlight SourceNet’s progress. It substantially outperforms previous Southern California baselines (Ross et al., 2018; Cheng et al., 2023), which reported over $30^{\circ}$ deviations. While recent solvers like FOCONET (Song et al., 2025) and DiTing (Zhao et al., 2023) report lower errors than them ($20^{\circ}$$30^{\circ}$), they rely on dense networks or curated Quality A subsets. SourceNet matches this precision (Median $19^{\circ}$) on sparse, irregular operational data. Also, this performance lies within the $20^{\circ}$$30^{\circ}$ discrepancy range typically observed between different high-quality manual catalogs, indicating that SourceNet has bridged automated inference with human-analyst reliability under real-world conditions.
5.3 Ablation Studies
To validate our architecture, we perform component-wise ablation studies on the synthetic test set. We train two variants: one removing the Scalar Tower (waveform only), and one replacing the Set Transformer with a DeepSets-style pooling architecture. The results are summarized in Table 1. We found that removing the scalar metadata causes the Magnitude prediction to collapse as MAE increasing by over 150% (0.07 $\to$ 0.18). This shows that waveform shapes alone are ambiguous on energy scale, and explicit geometric context (distance, attenuation) fed by the Scalar Tower is essential for disentangling energy from propagation effects. Also, we found replacing the Transformer with simple pooling (DeepSets) makes the mechanism estimation worse: the $M_{ij}$ error nearly triples (0.06 $\to$ 0.17), and the Kagan angle degrades significantly ($10.2^{\circ}\to 28.4^{\circ}$). This suggests that wave physics is relational: determining a source mechanism requires comparing relative data information across the array. Self-Attention captures these pairwise dependencies, whereas global pooling masks them.
Table 1: Ablation Study. Impact of architectural components on model performance. Removing the Scalar Tower degrades Magnitude prediction, while removing Self-Attention (replacing it with pooling) severely degrades the Moment Tensor mechanism estimation.
Model Variant
Mag ($M_{w}$)
MAE
Tensor ($M_{ij}$)
Mean MAE
Kagan Angle
Mean ()
SourceNet (Full) 0.07 0.06 10.2
w/o Scalar Tower 0.18 0.09 12.5
w/o Self-Attention 0.10 0.17 28.4
6 Interpretability: Emergent Physical Strategies
Deep learning models are often criticized as black boxes (Rudin, 2019). Here, we show that SourceNet is not just fitting function, but has autonomously discovered physical strategies that extend human intuitions (Iten et al., 2020).
A question of interest is how much of a waveform is actually useful. Classical solvers typically rely on a single data point: the first P-wave polarity. The rest of the time series is often discarded as coda noise due to complex scattering. However, our Grad-CAM analysis (Figure 5) shows that SourceNet’s attention is not confined to the P-wave onset but extends substantially into the coda and early scattered phases. Physically, this suggests the model has learned that these phases are not random but contains rich information about the source mechanism (Aki & Richards, 2002). Thus, SourceNet effectively performs an implicit full-waveform inversion from the entire signal duration that human analysts usually ignore (Fichtner, 2010).
The most interesting emergent behavior appears in the Transformer’s self-attention mechanism. When aggregating attention scores by station azimuth (Figure 5), we observe a systematic anisotropy: the model consistently upweights stations in the East-West direction ($90^{\circ}/270^{\circ}$) and downweights the denser stations to the North and South. This attention bias is not an artifact of specific source parameter preference (e.g., the prevalence of strike-slip faulting) in the real data, because we observed the same pattern in the pre-trained model on synthetic data with uniformly distributed source mechanisms. Thus, the learned policy is driven by the network topology itself. The model has successfully decoupled the sensor geometry from the wave physics, and identified the intrinsic spatial limitations of the sensor array.
Figure 5: Emergent Scientific Discovery: Learning How to Observe. (a) Learned local strategy. Grad-CAM visualization reveals that the model attends significantly to the P-coda and early scattered phases (red regions), effectively performing an implicit full-waveform inversion using data traditionally discarded as noise. (b) Learned global policy. Aggregated attention weights (bar plot) reveal a learned anisotropy, prioritizing East-West stations. As shown in the station map (c), this directly counters the network’s geometric bias (dense N-S coverage along the fault). The model autonomously identified the information bottleneck and recovered principles of Optimal Experimental Design by prioritizing under-sampled orthogonal views.
This behavior represents an autonomous discovery of information geometry. The dense N-S sensors, while numerous, are highly correlated in terms of the constraints they provide, and adding more yields diminishing returns. SourceNet has identified this information bottleneck (Tishby et al., 2000). Its attention mechanism acts as a dynamic preconditioner by performing a soft selection that downweights redundant observations to mitigate multicollinearity, while prioritizing the under-sampled orthogonal views (E-W) to stabilize the inversion. This demonstrates that the model has implicitly found the D-Optimality design (Pukelsheim, 2006; Krause et al., 2008), i.e. maximizing the determinant of the Fisher Information Matrix by focusing on the most variance-reducing observations.
7 Theoretical Implications
Amortized Inference vs. Classical Optimization. Classical geophysical inversion treats every event as a new optimization puzzle ($O(K)$), which makes it sensitive to initialization and computationally expensive. SourceNet represents a paradigm shift towards amortized inference (Gershman & Goodman, 2014). By paying the heavy computational cost upfront during training, the model learns the global inverse operator to allow instantaneous ($O(1)$) prediction (Adler & Öktem, 2017; Ongie et al., 2020). Moreover, unlike deterministic solvers that fit a single solution, the model learns a posterior mean that naturally marginalizes out the random uncertainties (e.g. velocity model errors) introduced during training, which leads to robustness by design.
Self-Attention as Data-Driven Green’s Functions. Mathematically, the Set Transformer implements a form of learned kernel regression (Tsai et al., 2019), so that its self-attention mechanism computes a similarity matrix that somewhat acts as a data-driven Green’s function (Kovachki et al., 2023). Instead of relying on analytical wave equations, which can be too ideal, the model learns to weigh sensor contributions based on their signal data and geometric leverage. So physically, it perform beamforming by dynamically propagating information on the observations that maximize information gain.
PSDR as Causal Representation Learning. From a causal inference perspective, our Physics-Structured Domain Randomization is not simply data augmentation, but an intervention on nuisance variables (Pearl, 2009). By randomizing the propagation environment (velocity models, noise, topology) while keeping the source parameters fixed, we mathematically force the encoder to discard spurious correlations (e.g., path and site effects) and learn source physics, which is the invariant representation between simulation and reality (Arjovsky et al., 2019; Schölkopf et al., 2021).
Limitations and Future Work. Our current framework assumes that the support of the simulation covers the physics of the real world. Out-of-distribution (OOD) events, such as complex multi-fault ruptures are not modeled in our synthetics and remain challenging. Looking forward, we will extend our framework to model the full spatiotemporal evolution of the fault slip using neural surrogates (Li et al., 2020; Karniadakis et al., 2021), to constrain the dynamic physics directly from observational data. This is a move from static parameter estimation to high-resolution dynamic process characterization.
8 Conclusion
In this work, we proposed SourceNet, a framework that bridges deep learning and rigorous physical inference on variable-geometry arrays. By treating sensor arrays as flexible sets and bridging the Sim-to-Real gap through physics-structured randomization, we demonstrated that AI can resolve the non-linear complexities of wave propagation without relying on massive labeled real-world datasets. Our findings offer two broader insights for the AI for Science community. First, architecture should follow physics. The success of our Set Transformer over pooling baselines confirms that capturing relational dependencies (e.g., relative phases) is theoretically needed for wavefield inversion. Second, interpretability leads to discovery. SourceNet proved to be more than a predictor, as it acted as a scientific agent as it autonomously discovers the geometry gap of the sensor network and recovers optimal experimental design from data alone. By enabling models to reason over sparse, irregular observations, we address a bottleneck in scaling AI to the physical world. Towards physical Foundation Models, the ability to learn invariant operators from randomized physics, and to diagnose the limitations of the observing system itself, will be critical for turning data to scientific discovery.